OpenStudy (anonymous):

System of equation: 2.5x - 3y = -13 3.25x - y = -14 I got to 7.25x = 27; x = 3.7241... Am I correct?

OpenStudy (anonymous):

My full work: y = mx + c form: 3.25x − y = −14 y = 3.25x + 14 Input y into the first equation: 2.5x − 3y = −13 2.5x − 3(3.25x+14) = −13 2.5x − 9.75x = −13 − 42 −7.25x = −55 7.25x = 27

OpenStudy (whpalmer4):

\[3.25x - y = -14\] \[y = 3.25x + 14\] \[2.5x - 3y = -13\]\[2.5x-3(3.25x+14) = -13\]\[2.5x-9.75x-42=-13\]\[-7.25x=-13+42\] try again from there...

OpenStudy (anonymous):

Thanks again :)

OpenStudy (whpalmer4):

what is your final value for x and y?

OpenStudy (anonymous):

x = -4 y = 1

OpenStudy (whpalmer4):

Good job!